OpenStudy (sodapop):

Simplify (1- cos θ)(1+cos θ) /(1-sin θ)(1+sin θ)

hero (hero):

You are given \[\frac{(1 - \cos(\theta))(1 + \cos(\theta))}{(1 - \sin(\theta))(1 + \sin(\theta))}\] and asked to simplify the expression right?

OpenStudy (sodapop):

Correct

OpenStudy (sodapop):

\[(1- \cos θ)(1+\cos θ) \over (1-\sin θ)(1+\sin θ)\]

hero (hero):

And in this case can apply the difference of squares rule \((a - b)(a + b) = a^2 - b^2\) so that: \(\dfrac{(1 - \cos(\theta))(1 + \cos(\theta))}{(1 - \sin(\theta))(1 + \sin(\theta))} = \dfrac{1 - \cos^2(\theta)}{1 - \sin^2(\theta)}\) Do you agree?

OpenStudy (sodapop):

Yes I do

hero (hero):

Now from Pythagorean Identity \(\sin^2\theta + \cos^2 \theta = 1\), we know that \(\sin^2 \theta = 1 - \cos^2\theta\) and \(\cos^2\theta = 1 - \sin^2\theta\) right?

OpenStudy (sodapop):

Yeah

hero (hero):

Therefore \[\dfrac{1 - \cos^2(\theta)}{1 - \sin^2(\theta)} = \dfrac{\sin^2\theta}{\cos^2\theta}\] Correct?

OpenStudy (sodapop):

Yeah thank you so much for explaining

hero (hero):

We're not done yet. Because of algebraic rule \(\dfrac{a^2}{b^2} = \left(\dfrac{a}{b}\right)^2\) We can write \(\dfrac{\sin^2\theta}{\cos^2\theta} = \left(\dfrac{\sin\theta}{\cos\theta}\right)^2\). Do you agree?

OpenStudy (sodapop):

I do agree

hero (hero):

And we know that \(\tan \theta = \dfrac{\sin \theta}{\cos \theta}\) therefore \(\left(\dfrac{\sin\theta}{\cos\theta}\right)^2 = (\tan\theta)^2\)

hero (hero):

Makes sense right?

OpenStudy (sodapop):

Yes it does

hero (hero):

So in conclusion should we should end by stating \(\dfrac{(1 - \cos(\theta))(1 + \cos(\theta))}{(1 - \sin(\theta))(1 + \sin(\theta))} = \tan^2\theta\) ?

OpenStudy (sodapop):

Well thank you Hero for helping. I understand how to do it now.

OpenStudy (sodapop):

You were a big help

hero (hero):

You're welcome